# On the fixed locus of the antisymplectic involution of an EPW cube

### v1

@@ 4.1. The special fiber of $\mathscr{X}\to D$
which is the identity on $X\smallsetminus\pi(\Delta)\simeq M_{\sigma}(v)\smallsetminus\Delta$ to $[F]$, and that sends $[F]\in M_{S}(a)\simeq\pi(\Delta)$ to the polystable object $[\mathscr{O}_{S}(H)\oplus F]$.
@@ 4.2. Fixed locus on the special fiber
  • ${F_{2}}$ is irreducible of dimension 2 and contained in the fourfold $W=\pi(\Delta)$;
@@
$\operatorname{Fix}(j^{[3]})=\Gamma^{[3]}\sqcup T.$
Note that $T$ is contained in the exceptional divisor $\Delta$. Indeed, if $Z\in S^{[3]}$ is fixed by $j^{[3]}$, then its support $\operatorname{Supp}(Z)$ is either contained in $\Gamma$, or of the form $\operatorname{Supp}(Z)=\{x+j(x)+y\}$. In particular, if $Z\notin\Gamma^{[3]}$ is fixed by $j^{[3]}$, then $\varphi(Z)$ is supported on two points, hence lies on a line in ${\bf P}^{2}$, so that $Z\in\Delta$.
This characterizes the fixed locus of $\iota$ outside the smooth fourfold $W=\pi(\Delta)\subseteq X$. We study the action of $\iota$ on $W=M_{S}(0,L,-4)$.
@@
The action of the involution $\iota$ on the smooth fourfold $W=M_{S}(0,L,-4)$ is given by
@@
for any $C\in|L|$ and any degree $-3$ torsion-free rank 1 sheaf $\xi_{3}$ on $C$.
@@
Suppose that $C$ is smooth and $\xi_{3}=\mathscr{O}_{C}(-Z)$, where $Z=p+q+r\in C^{(3)}$. Then it satisfies $\iota\left(i_{C,*}\left(\mathscr{O}_{C}(-Z)\right)\right)=i_{C,*}\left(\mathscr{O}_{C}\left(-j^{[3]}(Z)\right)\right)$.
By the adjunction formula, the canonical bundle of $C$ is equal to $\mathscr{O}_{C}(L)$. Since $\varphi|_{C}\colon C\to{\bf P}^{1}$ is a double cover, for any point $p\in C$, we have isomorphisms
@@
In particular, formula (19) is verified for any element of $W$ of the form $i_{C,*}(\mathscr{O}_{C}(-p-q-r))$ where $p,q,r$ are points on $C$ and $C\in|H|$ is a smooth curve.
By Bertini’s theorem, these elements form a dense open subset of $W$, therefore the involution on $W$ is given by formula (19) everywhere. ∎
@@
Under this isomorphism, the involution $\iota$ on $W$ corresponds to the involution $\iota_{0}$ on $M_{S}(0,L,0)$ given by
@@
which is one of the involutions studied in [FMOS1].
Using the description of the fixed locus of $\iota_{0}$ in [FMOS1, Proposition 4.2], we obtain that the fixed locus of $\iota|_{W}$ has two connected components $W_{+}$ and $W_{-}$, both smooth irreducible and 2-dimensional, which are the closures of the loci of $i_{C,*}(\xi\otimes\mathscr{O}(-L))$ where $\xi$ is an even, or odd, theta characteristic on a smooth curve $C$. Since $C$ has genus $2$, a theta characteristic $\xi$ on $C$ is odd if and only if $h^{0}(C,\xi)=1$ (the 6 odd theta characteristics correspond to the 6 Weierstrass points $x_{1},\dots,x_{6}$ of $C$; the 10 even theta characteristics correspond to the $x_{i}+x_{j}+x_{k}-g^{1}_{2}$, where $i,j,k\in\{1,\dots,6\}$ are distinct—this description counts each of them twice, using $x_{1}+\cdots+x_{6}=3g^{1}_{2}$).
To conclude, we show that $\operatorname{Fix}(\iota)$ is equal to the image $\pi(\operatorname{Fix}(j^{[3]}))$. Clearly, we have an inclusion $\pi(\operatorname{Fix}(j^{[3]}))\subseteq\operatorname{Fix}(\iota)$ and they are equal on $X\smallsetminus W$. We restrict ourselves to $W$, where $\operatorname{Fix}(\iota)\cap W=\operatorname{Fix}(\iota|_{W})=W_{+}\sqcup W_{-}$ and
$\pi\left(\operatorname{Fix}(j^{[3]})\right)\cap W=\pi\left(\operatorname{Fix}(j^{[3]})\cap\Delta\right)=\pi\left(\Gamma^{(3)}\cap\Delta\right)\cup\pi(T).$
Now, $\pi(T)$ is contained in $W_{+}\sqcup W_{-}$, which has dimension 2. This implies that $T$ is contracted by $\pi$. We show that $\pi(T)=W_{-}$. By definition of $W_{-}$, it is enough to show that for $i_{C,*}(\xi_{3})\in\pi(T)$ general, the element $\xi=\xi_{3}\otimes\mathscr{O}_{C}(L)$ is an odd theta characteristic. The general element of $\pi(T)$ is $\iota_{C,*}\mathscr{O}_{C}(-x-j(x)-y)\in W$, where $C$ is the pullback of the line through $\varphi(x)$ and $\varphi(y)$. Therefore, we obtain $\xi=\mathscr{O}_{C}(-x-j(x)-y)\otimes\mathscr{O}_{C}(x+j(x))=\mathscr{O}_{C}(-y)$, which is an odd theta characteristic by the discussion above, since $y$ belongs to $C\cap\Gamma$. Therefore, the morphism $\pi$ restricts to a ${\bf P}^{1}$-fibration $T\to W_{-}$.
The component $\Gamma^{(3)}$ intersects $\Delta$ in a divisor and is disjoint from $T$, hence $\pi(\Gamma^{(3)}\cap\Delta)$ is equal to $W_{+}$. The restriction of $\pi$ to $\Gamma^{3}\cap\Delta$ is finite, since the fiber over $i_{C,*}(\xi)$ is contained in $C^{(3)}\cap\Gamma^{(3)}\subseteq(C\cap\Gamma)^{3}$ Moreover, for any element $F=i_{C,*}(-Z)\in W_{+}$, for a smooth curve $C$, the intersection $\pi^{-1}(F)\cap\Gamma^{(3)}=|\mathscr{O}_{C}(-Z)|\cap\Gamma^{(3)}$ is given by $2$ points, namely $Z$ and the residual intersection $Z^{\prime}\coloneqq\Gamma\cdot C-Z$. Therefore, the induced morphism
[…]
@@
Therefore, $\pi\left(\operatorname{Fix}(j^{[3]})\right)\cap W=\operatorname{Fix}(\iota|_{W})$. In particular, the fixed locus of $\iota$ is the image of the fixed locus of $j^{[3]}$.
@@
The scheme $F_{3}$ is semi-log-canonical (see [K, Chapter 3, Definition 12]). Indeed, its normalization is the smooth scheme $\Gamma^{(3)}$ and the normalization morphism
@@ 4.3. Local description of $X$
Following [FMOS2], we give a local description of the variety $X$ and the involution $\iota$. Moreover, we show that the schematic fixed locus $\operatorname{Fix}(\iota)$ is a reduced local complete intersection.
@@
given by the Yoneda product. By [AS, Corollary 4.1], there is an isomorphism of germs
@@
The situation is analogous to that of [FMOS2, Remark 2.7 and Lemma 5.4]. Using their methods, we show the following lemma.
@@
Up to changing the singular point $[E]$ in its $S$-equivalence class, we can suppose that $E$ is of the form $T\oplus F$, where $T$ is the stable spherical sheaf $\mathscr{O}(-L)$ and $F$ is stable with Mukai vector $a=(0,L,-4)$. By Equation (21), the germ of $X$ at the singular point $E$ is equal to $\mu_{2}^{-1}(0)/\!\!/G$.
@@
Now the action of the group $G=\operatorname{Aut}(T)\oplus\operatorname{Aut}(F)\simeq({\bf C}^{*})^{2}$ is given by conjugation on elements of $\operatorname{Ext}^{1}(E,E)$. Therefore, it is trivial on $\operatorname{Ext}^{1}(F,F)$ and $(g,h)\in({\bf C}^{*})^{2}$ acts as the multiplication by $gh^{-1}$ on $\operatorname{Ext}^{1}(T,F)$ and as the multiplication by $hg^{-1}$ on $\operatorname{Ext}^{1}(F,T)$. Therefore, as in [FMOS2, Remark 2.7], the quotient $\mu_{2}^{-1}(0)/\!\!/G$ is equal to ${\bf A}_{\bf C}^{4}\times Q$, where
@@
Moreover the action of $\iota$ on $\operatorname{Ext}^{1}(F,T)$ is dual to the one on $\operatorname{Ext}^{1}(T,F)$ , therefore the action of $\iota$ on $Q$ is the one described in [FMOS2, Proposition 5.4]. In particular, its fixed locus is equal to $B=\operatorname{Spec}({\bf C}[u_{2},u_{3}]/u_{2}u_{3})$.
@@ 5. Degeneration of ${\mathsf{W}}_{A}$
Given the family $\mathscr{X}\to D$ studied in Section 4, we consider the schematic fixed locus $\operatorname{Fix}(\iota_{\mathscr{X}})\to\mathscr{D}$ of $\iota_{\mathscr{X}}$ on $\mathscr{X}$ (see the proof of [FMOS2, Proposition 5.6] for the definition of the schematic structure on the fixed locus). For $t\neq 0$, the fiber $\operatorname{Fix}(\iota_{\mathscr{X}})_{t}=\operatorname{Fix}(\iota_{t})$ is the smooth Lagrangian threefold ${\mathsf{W}}_{A_{t}}$. The reducedness of the fixed locus of $\iota$ studied in Section 4.3 implies $\operatorname{Fix}(\iota_{\mathscr{X}})_{0}=\operatorname{Fix}(\iota)=F_{3}\cup F_{2}$.
@@
The morphism $f$ is flat by [L, 4.3 Proposition 3.9] since it is dominant and $\mathscr{Y}$ is integral. Therefore, it is equidimensional and $\mathscr{Y}_{0}$ has dimension $3$. Moreover, $\mathscr{Y}$ is contained in $\operatorname{Fix}(\iota_{\mathscr{X}})$, hence $\mathscr{Y}_{0}$ is equal to $F_{3}$ (scheme-theoretically).
@@
[…]
@@
All fibers of $f$ are slc: the central fiber is isomorphic to $F_{3}$, which is slc by Remark 4.4, and the other fibers ${\mathsf{W}}_{A}$ are smooth. Therefore, since $f$ is flat and proper, [K, Corollary 2.64] implies that $h^{p}(\mathscr{Y}_{t},\mathscr{O}_{\mathscr{Y}_{t}})$ is independent of $t\in D$. The theorem follows. ∎
@@ 5.1. Cohomology groups $H^{p}(F_{3},\mathscr{O}_{F_{3}})$
(25) 00 0 $\to H^{1}(F_{3},\mathscr{O}_{F_{3}})\to H^{1}(\Gamma^{(3)},\mathscr{O}_{\Gamma^{(3)}})\xrightarrow{\ \beta_{1}\ }H^{1}(W,\mathscr{R})\to H^{2}(F_{3},\mathscr{O}_{F_{3}})\to$
@@
(26) $H^{0}(\operatorname{Jac}(\Gamma),\Omega^{q}_{\operatorname{Jac}(\Gamma)})\longrightarrow H^{0}(\Gamma^{(3)},\Omega^{q}_{\Gamma^{(3)}})$
are isomorphisms for $q\in\{0,1,2,3\}$ ([M, (11.1)]). Therefore, it is enough to consider the restriction morphisms induced by the inclusion $E\subseteq\Gamma^{(3)}\subseteq\operatorname{Jac}(\Gamma)$. Moreover, since
(27) $H^{0}(\operatorname{Jac}(\Gamma),\Omega^{q}_{\operatorname{Jac}(\Gamma)})=\textstyle{\bigwedge\hskip-2.56073pt^{q}}\hskip 0.56905pt{H}^{0}(\operatorname{Jac}(\Gamma),\Omega^{1}_{\operatorname{Jac}(\Gamma)}),$
@@
We show that the surface $E\subseteq\operatorname{Jac}(\Gamma)$ is nondegenerate in the sense of [R, Section II], which implies by [R, Lemma II.1] that $r$ is injective.
Let $\theta$ be the canonical principal polarization on $\operatorname{Jac}(\Gamma)$. By [M, (16.2)], the cohomology class of $E$ in $H^{2}(\Gamma^{(3)},{\bf Z})$ is equal to
$-6\eta+\theta|_{\Gamma^{(3)}}\in H^{2}(\Gamma^{(3)},{\bf Z}),$
where $\eta\in H^{2}(\Gamma^{(3)},{\bf Z})$ is the class of $\Gamma^{(2)}\subseteq\Gamma^{(3)}$. [^4] The divisor $E\subseteq\Gamma^{(3)}$ is not ample. Indeed by [Ko, Lemma (1)], we have $(\theta-6\eta)^{3}=\sum_{i=0}^{3}{3\choose i}(-6)^{3-i}\frac{10!}{(10-i)!}=-36$ By the projection formula, the class of $E$ in $\operatorname{Jac}(\Gamma)$ is therefore given by
$[E]=-6[\Gamma^{(2)}]+[\Gamma^{(3)}]\cdot\theta.$
Since $[\Gamma^{(i)}]\in H^{*}(\operatorname{Jac}(\Gamma),{\bf Z})$ is equal to $\frac{\theta^{10-i}}{(10-i)!}$, it follows that $[E]=2\theta^{8}/8!$ in $H^{*}(\operatorname{Jac}(\Gamma),{\bf Z})$. Therefore, $[E]$ is a positive multiple of $\theta^{8}$, hence $E$ is nondegenerate by [R, Corollary II.2]. ∎

### v2

@@ 4.1. The special fiber of $\mathscr{X}\to D$
which is the identity on $X\smallsetminus\pi(\Delta)\simeq M_{\sigma}(v)\smallsetminus\Delta$ to $[F]$, and that sends $F]\in M_{S}(a)\simeq\pi(\Delta)$ to the polystable object $\mathscr{O}_{S}(H)\oplus F]$.
@@ 4.2. Fixed locus on the special fiber
  • $F_{2}$} is irreducible of dimension 2 and contained in the fourfold $T=\pi(\Delta)$;
@@
$\operatorname{Fix}(j^{[3]})=\Gamma^{[3]}\sqcup F^{\prime}.$
Note that $F^{\prime}$ is contained in the exceptional divisor $\Delta$. Indeed, if $Z\in S^{[3]}$ is fixed by $j^{[3]}$, then its support $\operatorname{Supp}(Z)$ is either contained in $\Gamma$, or of the form $\operatorname{Supp}(Z)=\{x+j(x)+y\}$. In particular, if $Z\notin\Gamma^{[3]}$ is fixed by $j^{[3]}$, then $\varphi(Z)$ is supported on two points, hence lies on a line in ${\bf P}^{2}$, so that $Z\in\Delta$.
This characterizes the fixed locus of $\iota$ outside the smooth fourfold $W=\pi(\Delta)\subseteq X$. We study the action of $\iota$ on $T=M_{S}(0,L,-4)$.
@@
The action of the involution $\iota$ on the smooth fourfold $T=M_{S}(0,L,-4)$ is given by
@@
for any $C\in|L|$ and any degree 3-3 torsion-free rank 1 sheaf $\xi_{3}$ on $C$.
@@
Suppose that $C$ is smooth and $\xi_{3}=\mathscr{O}_{C}(-Z)$, where $Z=p+q+r\in C^{(3)}$. Then it satisfies $\iota\left(i_{C,*}\left(\mathscr{O}_{C}(-Z)\right)\right)=i_{C,*}\left(\mathscr{O}_{C}\left(-j^{[3]}(Z)\right)\right)$. By the adjunction formula, the canonical bundle of $C$ is equal to $\mathscr{O}_{C}(L)$. Since $\varphi|_{C}\colon C\to{\bf P}^{1}$ is a double cover, for any point $p\in C$, we have isomorphisms
@@
In particular, formula (21) is verified for any element of $T$ of the form $i_{C,*}(\mathscr{O}_{C}(-p-q-r))$ where $p,q,r$ are points on $C$ and $C\in|H|$ is a smooth curve.
By Bertini’s theorem, these elements form a dense open subset of $T$, therefore the involution on $T$ is given by formula (21) everywhere. ∎
@@
Under this isomorphism, the involution $\iota$ on $T$ corresponds to the involution $\iota_{0}$ on $M_{S}(0,L,0)$ given by
@@
which is one of the involutions studied in [FMOS22].
Using the description of the fixed locus of $\iota_{0}$ in [FMOS22, Proposition 4.2], we obtain that the fixed locus of $\iota|_{T}$ has two connected components $W_{+}$ and $W_{-}$, both smooth irreducible and 2-dimensional, which are the closures of the loci of $i_{C,*}(\xi\otimes\mathscr{O}(-L))$ where $\xi$ is an even, or odd, theta characteristic on a smooth curve $C$. Since $C$ has genus $2$, a theta characteristic $\xi$ on $C$ is odd if and only if $h^{0}(C,\xi)=1$ (the 6 odd theta characteristics correspond to the 6 Weierstrass points $x_{1},\dots,x_{6}$ of $C$; the 10 even theta characteristics correspond to the $x_{i}+x_{j}+x_{k}-g^{1}_{2}$, where $i,j,k\in\{1,\dots,6\}$ are distinct—this description counts each of them twice, using $x_{1}+\cdots+x_{6}=3g^{1}_{2}$).
To conclude, we show that $\operatorname{Fix}(\iota)$ is equal to the image $\pi(\operatorname{Fix}(j^{[3]}))$. Clearly, we have an inclusion $\pi(\operatorname{Fix}(j^{[3]}))\subseteq\operatorname{Fix}(\iota)$ and they are equal on $X\smallsetminus T$. We restrict ourselves to $T$, where $\operatorname{Fix}(\iota)\cap T=\operatorname{Fix}(\iota|_{T})=W_{+}\sqcup W_{-}$ and
$\pi\left(\operatorname{Fix}(j^{[3]})\right)\cap W=\pi\left(\operatorname{Fix}(j^{[3]})\cap\Delta\right)=\pi\left(\Gamma^{(3)}\cap\Delta\right)\cup\pi(F^{\prime}).$
Now, $\pi(F^{\prime})$ is contained in $W_{+}\sqcup W_{-}$, which has dimension 2. This implies that $F^{\prime}$ is contracted by $\pi$. We show that $\pi(F^{\prime})=W_{-}$. By definition of $W_{-}$, it is enough to show that for $i_{C,*}(\xi_{3})\in\pi(F^{\prime})$ general, the element $\xi=\xi_{3}\otimes\mathscr{O}_{C}(L)$ is an odd theta characteristic. The general element of $\pi(F^{\prime})$ is $\iota_{C,*}\mathscr{O}_{C}(-x-j(x)-y)\in T$, where $C$ is the pullback of the line through $\varphi(x)$ and $\varphi(y)$. Therefore, we obtain $\xi=\mathscr{O}_{C}(-x-j(x)-y)\otimes\mathscr{O}_{C}(x+j(x))=\mathscr{O}_{C}(-y)$, which is an odd theta characteristic by the discussion above, since $y$ belongs to $C\cap\Gamma$. Therefore, the morphism $\pi$ restricts to a ${\bf P}^{1}$-fibration $F^{\prime}\to W_{-}$.
The component $\Gamma^{(3)}$ intersects $\Delta$ in a divisor and is disjoint from $F^{\prime}$, hence $\pi(\Gamma^{(3)}\cap\Delta)$ is equal to $W_{+}$. The restriction of $\pi$ to $\Gamma^{3}\cap\Delta$ is finite, since the fiber over $i_{C,*}(\xi)$ is contained in $C^{(3)}\cap\Gamma^{(3)}\subseteq(C\cap\Gamma)^{3}$ Moreover, for any element $F=i_{C,*}(-Z)\in W_{+}$, for a smooth curve $C$, the intersection $\pi^{-1}(F)\cap\Gamma^{(3)}=|\mathscr{O}_{C}(-Z)|\cap\Gamma^{(3)}$ is given by $2$ points, namely $Z$ and the residual intersection $Z^{\prime}\coloneqq\Gamma\cdot C-Z$. Therefore, the induced morphism
(22) $\Gamma^{(3)}\cap\Delta\xrightarrow{\ 2:1\ }W_{+}$
@@
Therefore, $\pi\left(\operatorname{Fix}(j^{[3]})\right)\cap T=\operatorname{Fix}(\iota|_{T})$. In particular, the fixed locus of $\iota$ is the image of the fixed locus of $j^{[3]}$.
@@
The scheme $F_{3}$ is semi-log-canonical (see [Kol23, Chapter 3, Definition 12]). Indeed, its normalization is the smooth scheme $\Gamma^{(3)}$ and the normalization morphism
@@ 4.3. Local description of $X$
Following [FMOS23], we give a local description of the variety $X$ and the involution $\iota$. Moreover, we show that the schematic fixed locus $\operatorname{Fix}(\iota)$ is a reduced local complete intersection.
@@
given by the Yoneda product. By [AS25, Corollary 4.1], there is an isomorphism of germs
@@
The situation is analogous to that of [FMOS23, Remark 2.7 and Lemma 5.4]. Using their methods, we show the following lemma.
@@
Up to changing the singular point $[E]$ in its $S$-equivalence class, we can suppose that $E$ is of the form $T\oplus F$, where $T$ is the stable spherical sheaf $\mathscr{O}(-L)$ and $F$ is stable with Mukai vector a=(0,L,4)a=(0,L,-4). By Equation (23), the germ of $X$ at the singular point $E$ is equal to $\mu_{2}^{-1}(0)/\!\!/G$.
@@
Now the action of the group $G=\operatorname{Aut}(T)\oplus\operatorname{Aut}(F)\simeq({\bf C}^{*})^{2}$ is given by conjugation on elements of $\operatorname{Ext}^{1}(E,E)$. Therefore, it is trivial on $\operatorname{Ext}^{1}(F,F)$ and $(g,h)\in({\bf C}^{*})^{2}$ acts as the multiplication by $gh^{-1}$ on $\operatorname{Ext}^{1}(T,F)$ and as the multiplication by $hg^{-1}$ on $\operatorname{Ext}^{1}(F,T)$. Therefore, as in [FMOS23, Remark 2.7], the quotient $\mu_{2}^{-1}(0)/\!\!/G$ is equal to ${\bf A}_{\bf C}^{4}\times Q$, where
@@
Moreover the action of $\iota$ on $\operatorname{Ext}^{1}(F,T)$ is dual to the one on $\operatorname{Ext}^{1}(T,F)$, therefore the action of $\iota$ on $Q$ is the one described in [FMOS23, Proposition 5.4]. In particular, its fixed locus is equal to $B=\operatorname{Spec}({\bf C}[u_{2},u_{3}]/u_{2}u_{3})$.
@@ 5. Degeneration of ${\mathsf{W}}_{A}$
Given the family $\mathscr{X}\to D$ studied in Section 4, we consider the schematic fixed locus $\operatorname{Fix}(\iota_{\mathscr{X}})\to\mathscr{D}$ of $\iota_{\mathscr{X}}$ on $\mathscr{X}$ (see the proof of [FMOS23, Proposition 5.6] for the definition of the schematic structure on the fixed locus). For $t\neq 0$, the fiber $\operatorname{Fix}(\iota_{\mathscr{X}})_{t}=\operatorname{Fix}(\iota_{t})$ is the smooth Lagrangian threefold ${\mathsf{W}}_{A_{t}}$. The reducedness of the fixed locus of $\iota$ studied in Section 4.3 implies $\operatorname{Fix}(\iota_{\mathscr{X}})_{0}=\operatorname{Fix}(\iota)=F_{3}\cup F_{2}$.
@@
The morphism $f$ is flat by [Liu02, 4.3 Proposition 3.9] since it is dominant and $\mathscr{Y}$ is integral. Therefore, it is equidimensional and $\mathscr{Y}_{0}$ has dimension $3$. Moreover, $\mathscr{Y}$ is contained in $\operatorname{Fix}(\iota_{\mathscr{X}})$, hence $\mathscr{Y}_{0}$ is equal to $F_{3}$ (scheme-theoretically).
@@
Let $A\in\operatorname{\mathsf{LG}}(\textstyle{\bigwedge\thinspace^{3}}\thinspace{V}_{6})\smallsetminus\Sigma\cup\Gamma$ be a Lagrangian, with associated smooth EPW cube $\widetilde{\mathsf{Z}}_{A}$. The external Hodge numbers of ${\mathsf{W}}_{A}$ can be computed as
@@
All fibers of $f$ are slc: the central fiber is isomorphic to $F_{3}$, which is slc by Remark 4.4, and the other fibers ${\mathsf{W}}_{A}$ are smooth. Therefore, since $f$ is flat and proper, [Kol23, Corollary 2.64] implies that $h^{p}(\mathscr{Y}_{t},\mathscr{O}_{\mathscr{Y}_{t}})$ is independent of $t\in D$. The theorem follows. ∎
@@ 5.1. Cohomology groups $H^{p}(F_{3},\mathscr{O}_{F_{3}})$
(27) $0$ $\to H^{1}(F_{3},\mathscr{O}_{F_{3}})\to H^{1}(\Gamma^{(3)},\mathscr{O}_{\Gamma^{(3)}})\xrightarrow{\ \beta_{1}\ }H^{1}(W,\mathscr{R})\to H^{2}(F_{3},\mathscr{O}_{F_{3}})\to$
@@
[…]
are isomorphisms for $q\in\{0,1,2,3\}$ ([Mac62, (11.1)]). Therefore, it is enough to consider the restriction morphisms induced by the inclusion $E\subseteq\Gamma^{(3)}\subseteq\operatorname{Jac}(\Gamma)$. Moreover, since
[…]
@@
We show that the surface $E\subseteq\operatorname{Jac}(\Gamma)$ is nondegenerate in the sense of [Ran80, Section II], which implies by [Ran80, Lemma II.1] that $r$ is injective.
Let $\theta$ be the canonical principal polarization on $\operatorname{Jac}(\Gamma)$. By [Mac62, (16.2)], the cohomology class of $E$ in $H^{2}(\Gamma^{(3)},{\bf Z})$ is equal to
$6\eta+\theta|_{\Gamma^{(3)}}\in H^{2}(\Gamma^{(3)},{\bf Z}),$
where $\eta\in H^{2}(\Gamma^{(3)},{\bf Z})$ is the class of $\Gamma^{(2)}\subseteq\Gamma^{(3)}$. [^4] The divisor $E\subseteq\Gamma^{(3)}$ is not ample. Indeed by [Kou93, Lemma (1)], we have $(\theta-6\eta)^{3}=\sum_{i=0}^{3}{3\choose i}(-6)^{3-i}\frac{10!}{(10-i)!}=-36$ By the projection formula, the class of $E$ in $\operatorname{Jac}(\Gamma)$ is therefore given by
$E]=-6[\Gamma^{(2)}]+[\Gamma^{(3)}]\cdot\theta.$
Since $\Gamma^{(i)}]\in H^{*}(\operatorname{Jac}(\Gamma),{\bf Z})$ is equal to $\frac{\theta^{10-i}}{(10-i)!}$, it follows that $E]=2\theta^{8}/8!$ in $H^{*}(\operatorname{Jac}(\Gamma),{\bf Z})$. Therefore, $[E]$ is a positive multiple of $\theta^{8}$, hence $E$ is nondegenerate by [Ran80, Corollary II.2]. ∎